OpenStudy (anonymous):

Can someone tell me the answer to this if I'm trying to turn this into slope intercept form? 3/4y -6x=3

OpenStudy (aravindg):

well slope intercept form is y-mx+c so rearrange this equation in that form !

OpenStudy (phi):

as a first step, try adding +6x to both sides. what do you get?

OpenStudy (anonymous):

3/4 y = 6x+3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

looks good, now multiply both sides by 4/3 that means multiply every term on both sids by 4/3 of course, we expect 4/3 *3/4 to be 1, so you get y = on the left side what do you get for the entire equation ?

OpenStudy (anonymous):

y= 6x+3 @phi

OpenStudy (phi):

you start with 3/4 y = 6x+3 multiply BOTH sides by 4/3 (all terms)

OpenStudy (anonymous):

OH! I got it! thanks for your help!! @phi

OpenStudy (phi):

what do you get ?